OpenStudy (anonymous):

What is the sum of a 22-term arithmetic sequence where the first term is 57 and the last term is -27? What is the sum of a 6-term geometric series if the first term is 24 and the last term is 1,417,176?

OpenStudy (campbell_st):

claire can you write this formula down sum of an arithmetic series given 1st and last term \[s _{n} = \frac{n}{2}(a + l)\] n = number of terms , a = 1st term l = last term substitute your information and evaluate

OpenStudy (anonymous):

I was not given that formula, I only have a1-a1r^n/1-r. Thank you!

OpenStudy (campbell_st):

Part 2 a term in geometric series is found using \[T _{n} = ar^{n-1}\] a = 1st term, r = common ratio and n = number of terms \[1417176=24r^5\] solve for r divide by 24 \[r^5 = 59049\] take the 5th root r = 9 now use the sum of a geometric series \[s _{n} = \frac{a(r^n - 1)}{r -1}\] in your question, a = 24, r = 9 and n = 6 good luck